Please help me with this question, thank you very much!
Q = 30L-3L2 + 60K-3K2
1. Find the cost minimizing amount of L and K to produce 60 units of output when w=1 and r=2.
2. Using the results from question 1, if the price of Q is $10/unit, calculate your profits froproducing 60 units.
Answer #1

1. By using solver in excel we get results,

Q

L

K

TC

60

0.42

0.83

125.23

TC = 60*(L*1+K*2), minimizing TC, by changing L. K and keeping Q=60

2.  

Q

L

K

TC

Price

TR

Profit

60

0.42

0.83

125.23

10

600.00

474.77

Profit = TR-TC

TR=60*10